OpenStudy (anonymous):

One of the largest rectangular omelets ever cooked had a perimeter of 80 feet. The length was 10 feet more than twice its width. How many feet wide was the omelet?

OpenStudy (anonymous):

l= w + 10 l = w = 80

OpenStudy (anonymous):

l + w= 80*

OpenStudy (anonymous):

i guess it might be a rectangle hence as said let width be; x, hence length will be ; 2x + 10 perimeter is 2l + 2w hence add substituting the values 2(x + 2x +10) = 80 divide by 2 3x + 10 = 40 subtract be 10 3x = 30 x = 10 hence width and length are 30 and 70

OpenStudy (anonymous):

But that together is already more thatn 80...

OpenStudy (anonymous):

trueeee^^^

OpenStudy (anonymous):

oops i wrote x = 10 m such a dumb; length = 10 and width equals 30

OpenStudy (anonymous):

i guess it might be a rectangle hence as said let width be; x, hence length will be ; 2x + 10 perimeter is 2l + 2w hence add substituting the values 2(x + 2x +10) = 80 divide by 2 3x + 10 = 40 subtract be 10 3x = 30 x = 10 hence width and length are 10 and 30 respectively, heres whole solution ;)